OpenStudy (anonymous):

(3x^4y^8/6xy^2)^-4 simplify

jhonyy9 (jhonyy9):

3x^4y^8) x^3y^6 2 16 (---------)^(-4) =( --------- )^(-4) = ( ---------- )^4 = ----------- 6xy^2 2 x^3y^6 x^12y^24